Which of the following equations are equivalent to -2m - 5m - 8 = 3 + (-7) + m?
-15m = -4m
-7m - 8 = m - 4
-3m - 8 = 4 - m
m - 4 = -7m - 8
-8 - 7m = -4 + m
-8 - 3m = 4 - m

Answer by unlockmath(1688)

Hello,
This -2m - 5m - 8 = 3 + (-7) + m can be rewritten as:
-7m - 8 = -4 + m
You should be able to figure out from this which ones are the same.
RJ
www.math-unlock.com